Marquis Company uses a weighted-average perpetual inventory system.

August 2, 22 units were purchased at $3 per unit.
August 18, 27 units were purchased at $5 per unit.
August 29, 24 units were sold.


What is the amount of the cost of goods sold for this sale?
Answer #1

Solution,

22 units * $3per unit = $66

27 units * $5 per unit = $135

Weighted average cost pr unit = (66+135)/49 = $4.102 per unit

Therefore, Cost of goods sold = $4.102 * 24 = $98.449